OpenStudy (anonymous):

(x+1)/(x-5) divide (x-2)/ (x^2-7x+10)

sam (.sam.):

\[\frac{(x+1)}{(x-5)} \div \frac{(x-2)}{(x^2-7x+10)}\] Switch the divide to multiply \[\frac{(x+1)}{(x-5)} \times \frac{(x^2-7x+10)}{(x-2)}\] Then multiply them together

OpenStudy (anonymous):

do i multiply straight across? or cross multiply

sam (.sam.):

straight across, you cross multiply only when there's '=' sign in between

OpenStudy (anonymous):

ohhhh i gotcha thanks :D